# ARIM-mdx Data System: Towards a Nationwide Data Platform for Materials Science

### v1

@@ V-B General Statistic
We present the statistical data from the one-year trial operation (from Aug. 2023 to Aug. 2024). Table I shows the overall information for the annual usage. The number of activated users are 837 from over 140 organizations, where 73.7 % of the users are from academia while 26.3 % are from industry. There are 377 research projects utilizing the ARIM-mdx data system, with the average of approximately 2.48 researchers per project. The total data volume is approximately 40.6 TB, where 11% of them are experimental data while 89% are theoretical data. The total number of data files is approximately 9.05 millions, with more experimental data files than theoretical ones, in contrast to the data volume distribution. There are 50 data transfer IoT devices and 6 connected supercomputers. We have been making steady progress in establishing the connections between the ARIM-mdx data system and facilities/supercomputers.
@@
# of Facilities Connected to Data Transfer IoT devices 50
# of Supercomputers Utilized by ARIM-mdx’s Users 6
Fig. 99 shows the increase in the total number of users, the volume of data, and the number of data files in the system, respectively. The number of users increased by 12.7 times over the course of a year (from Aug. 2023 to Aug. 2024). The number increases dramatically in May because we intensively conducted tutorial sessions. The significant increases in the data volume and file numbers from March to April are due to intensive physical simulations in one theoretical research project.
The overall results show that we successfully attracted a significant number of users during the one-year trial period. In particular, we successfully garnered significant interest from not only researchers on the academic organization but also the industrial organizations.
@@ VI Use Cases
In this section, we present several use cases on the ARIM-mdx data system. Since these use cases are ongoing, we describe only their key ideas, usage, and system aspects, without going into detail.
@@ VI-A Denoising Atomic Resolution Image
One of the primary goals of an atomic resolution electron microscopy is to enhance resolution, and the image denoising is a crucial method [51, 52]. Raw 2D image signals often contain information unrecognizable to the human eye due to inherent noise. By processing these raw images to a level where their embedded information becomes interpretable, denoising plays a vital role in supporting experiments, such as identifying optimal observation parts or parameters, and serves as a foundation for post-experimental theoretical analysis of physical phenomena.
Research tasks on the image denoising include both the interactive and batch workloads. For the interactive analysis, the denoising is executed for the target image data during the experiments as the examples for Jupyter Lab and Fuji/ImageJ, a popular image processing GUI tool in this field, on Jupyter remote desktop are shown in Fig. 13 and Fig. 13, respectively. We deploy the denoising program on the Jupyter environment of the ARIM-mdx data system in such a way that facility users immediately utilize the sophisticated denoising method on a many-core computing environment. The image data obtained from the facility computer is transferred via the IoT device. On the other hand, the batch processing is conducted on supercomputers for the post-experiment data analysis. In the post-experimental data analysis, the compute-intensive physical simulation, such as DFT calculation, is executed for building theoretical model of the target material. Moreover, for the development of a data-driven denoising method, supercomputers are utilized for the large-scale generation and training of the simulated image data. The trained model is deployed on the Jupyter side for the interactive inference.
Figure 10: Denoising for electron microscopy image. The configuration of this image is as follows: ADF-STEM CaF2 [001]; 512 $\times$ 512 pixels; accelerating voltage = 200 kV; convergence angle $\alpha$ = 22 mrad.
Figure 11: Jupyter Lab of ARIM-mdx data system.
Figure 12: Fiji / ImageJ on Jupyter Remote Desktop of ARIM-mdx data system.
Figure 13: VESTA in MateriApps Live! on Jupyter Remote Desktop of ARIM-mdx data system.
@@ VI-B Large-scale Physical Simulations
The large-scale theoretical data generation and collection for a physical simulation, such as DFT or MD, have been conducted on ARIM-mdx data system and supercomputers. The aim of the generation is, for example, to build a surrogate model for predicting physical characteristics or potentials, which are often time-consuming when calculating from scratch via the physical simulation. The key idea of the surrogate model is that given a large number of structure data for materials, we partly simulate the accurate results for the target characteristic, while the rest parts are trained by using a machine learning, such as graph neural networks (GNNs), and predicted. Since the data training is much faster than the one-by-one physical simulations, the surrogate model approach has a significant benefit on the efficiency.
The typical workflow of the data generation and collection includes the batch workloads on supercomputers and the interactive processing for the post-generation analysis. As shown in Fig. 13, the data collected on the ARIM-mdx data system can be interactively analyzed using GUI tools. MateriApps Live! installed on the Jupyter services includes most of the common tools (simulators, analyzers, and visualizers) for theoretical materials science, so that the developed programs on the ARIM-mdx data system can be immediately deployed and scaled up to supercomputers.
@@ VII Lessons and Future Directions
Over the course of a year of operation, we have obtained several insights and lessons. We summarize these insights and lessons together with the suggested future directions as follows:
Safe and efficient data extraction and storing from shared-use experimental facilities are very basic yet most crucial for all experimental researchers: According to the user’s survey, over 50% of users are only interested in the safe and efficient extraction of data, as typically each researcher already has the familiar tools for the data analysis. The access logs shows that the number of daily active users in Nextcloud is 10 times higher than in Jupyter services.
@@
Experimental data are inevitably much more sparse and smaller scale than theoretical data: Materials science experiments are inherently costly compared to computational methods and often require manual operations. The fact leads researchers to focus primarily on areas closely related to their target materials. Consequently, datasets generated by individual research projects tend to be localized within the broader landscape of materials. This situation presents a considerable challenge in the pursuit of developing more comprehensive and generalizable machine learning models or database. Specifically, to promote cross-material and general data utilization and publishing, addressing data sparsity has become an urgent priority.
For our future work, the ARIM-mdx data system plans to adopt multiple approaches. First, we aim to strengthen nationwide data collection efforts to build a more extensive materials’ database. Simultaneously, we will pursue integration with high-throughput automated experimental systems, enabling efficient and large-scale data generation. Furthermore, we plan to complement sparse data through the fusion of open data, academic literature, and state-of-the-art large language models (LLMs). Through these approaches, we aim to create a more comprehensive and versatile materials science database, supporting accelerated innovative materials development and scientific discoveries.
@@ VIII Related Work
Global efforts to promote open data [56] have led to the development of several data repository platforms for structuring and publishing scientific data, such as Zenodo [57], Dryad [58], and the others [59, 60, 61, 62]. The key difference from these general-purpose scientific data repositories is that ARIM-mdx data system focuses primarily on the collection and management of the raw data from the data source, rather than publishing the structured and cleaned data. The ARIM-mdx data system provides a data platform to effectively and efficiently prepare the structured dataset for the registration to these public repositories.
The integration of large-scale materials research facilities and world-class supercomputers has been undertaken at several national research institutes such as Oak Ridge National Laboratory, Argonne National Laboratory in the United States, and RIKEN in Japan, where both world-class supercomputer and large-scale material-related facilities are installed and connected [12, 13, 14, 15, 16, 17, 18, 19]. The key difference is that the ARIM-mdx data system focuses more on the nationwide deployment of the data system, where large-scale cross-institutional operations and infrastructure are crucial.
Nationwide and cross-institutional service operations for scientific research are well-established in high-performance computing fields. For example, in the United States, ACCESS [11] (formerly, XSEDE [9]) offers the shared-use supercomputing and storage services on 15 universities and research institutes. In the European Union, European Open Science Cloud (EOSC) [63] offers a federated and multidisciplinary computational environment for hosting and processing scientific research data. In Japan, the Japan High Performance Computing and Networking plus Large-scale Data Analyzing and Information Systems (JHPCN) [10] commonly offers the shared-use supercomputing services on 8 universities.
@@ IX Concluding Remarks
In this paper, we introduced the ARIM-mdx data system, a nationwide data platform for materials science in Japan. The ARIM-mdx data system is based on a dedicated high-performance storage system that is tightly integrated with supercomputers, the mdx cloud system, and experimental facilities across Japan via SINET6 and the data-transfer IoT devices. The ARIM-mdx data system achieves to efficiently store, manage, and analyze their data, facilitating more effective cross-organizational/domain research collaborations. As a result, the number of academic and industrial users approaches over 800 during the one-year trial operations. This highlights its potential as a reference design for large-scale modern materials-science data platforms.
Acknowledgement
The authors would like to thank Sohei Matsumoto, Takashi Endo, Yasukazu Murakami, Takahisa Yamamoto for the system deployment on AIST, Hokkaido University, Kyushu University, Nagoya University, and The University of Electro-Communications, respectively. The authors would also like to thank Taichi Kusumi for providing the experimental image data from the electron microscope.
This work was supported by ”Advanced Research Infrastructure for Materials and Nanotechnology in Japan (ARIM)” of MEXT, ”Joint Usage/Research Center for Interdisciplinary Large-scale Information Infrastructures (JHPCN)” in Japan (Project ID: jh241010), ”Developing a Research Data Ecosystem for the Promotion of Data-Driven Science” of MEXT, and JSPS KAKENHI Grant Number 22K17899.
We used mdx, supercomputer Wisteria/BDEC-01 at The University of Tokyo, Fugaku at RIKEN, Grand Chariot at Hokkaido University, MASAMUNE-IMR at Tokyoku University, Octpus at Osaka University, and ITO at Kyushu University.

### v2

@@ V-B General Statistic
We present the statistical data from the one-year operation (from Aug. 2023 to Aug. 2024). Table I shows the overall information for the annual usage. The number of activated users is 837 from over 140 organizations, where 73.7 % of the users are from academia while 26.3 % are from industry. There are 377 research projects utilizing the ARIM-mdx data system, with the average of approximately 2.48 researchers per project. The total data volume is approximately 40.6 TB, where 11% of them are experimental data while 89% are theoretical data. The total number of data files is approximately 9.05 million, with more experimental data files than theoretical ones, in contrast to the data volume distribution.
@@
Fig. 88 shows the increase in the total number of users, the volume of data, and the number of data files in the system, respectively. The number of users increased by 12.7 times over the course of a year (from Aug. 2023 to Aug. 2024). The number increases dramatically in May because we intensively conducted tutorial sessions. The significant increases in the data volume and file numbers in April are due to intensive physical simulations in one theoretical research project.
The overall results show that we successfully attracted a significant number of users during the one-year trial period. In particular, we successfully garnered significant interest from not only researchers in academic organizations but also industrial organizations.
Figure 9: Jupyter Lab of ARIM-mdx data system.
Figure 10: MateriApps LIVE! on Jupyter Remote Desktop.
@@ VI Use Cases
In this section, we present several use cases. As these are ongoing, we describe only their key ideas, usage, and system aspects, without details.
@@ VI-A Denoising Atomic Resolution Image
One of the primary goals of atomic resolution electron microscopy is to enhance resolution, and the image denoising is a crucial method [35, 36] as shown in Fig. 11. Raw 2D image signals often contain information unrecognizable to the human eye due to inherent noise. By processing these raw images to a level where their embedded information becomes interpretable, denoising plays a vital role in supporting experiments, such as identifying optimal observation parts or parameters, and serves as a foundation for post-experimental theoretical analysis of physical phenomena.
Research tasks in image denoising include both the interactive and batch workloads. For the interactive analysis, the denoising is executed for the target image data during the experiments as shown in Fig. 10. The image data from the facility computer is transferred via the IoT device. On the other hand, the batch processing is conducted on supercomputers for the post-experiment data analysis, where the physical simulation is executed for building the theoretical model.
Figure 11: Denoising for electron microscope image. The configuration of this image is as follows: ADF-STEM CaF2 [001]; 512 $\times$ 512 pixels; accelerating voltage = 200 kV; convergence angle $\alpha$ = 22 mrad [37].
@@ VI-B Large-scale Physical Simulations
The large-scale theoretical data generation and collection for a physical simulation, such as DFT or MD, have been conducted on ARIM-mdx data system and supercomputers. The typical workflow of the data generation and collection includes the batch workloads on supercomputers and the interactive processing for the post-generation analysis. As shown in Fig. 10 (using VESTA [38]), the data collected on the ARIM-mdx data system can be interactively analyzed using GUI tools. MateriApps LIVE! installed on the Jupyter services includes most of the common tools (simulators, analyzers, and visualizers) for theoretical materials science, so that the developed programs on the ARIM-mdx data system can be immediately deployed and scaled up to supercomputers.
@@ VII Lessons and Future Directions
Over the course of a year of operation, we have obtained several insights and lessons as follows:
Safe and efficient data extraction and storing from shared-use experimental facilities are very basic yet most crucial for all experimental researchers: According to the user’s survey, over 50% of users are only interested in the safe and efficient extraction of data, as typically each researcher already has the familiar tools for the data analysis. The access logs show that the number of daily active users in Nextcloud is 10 times higher than in Jupyter services.
@@
Experimental data are inevitably much more sparse and smaller scale than theoretical data: Materials science experiments are inherently costly compared to computational methods and often require manual operations. The fact leads researchers to focus primarily on areas closely related to their target materials. Consequently, datasets generated by individual research projects tend to be localized within the broader landscape of materials. This situation presents a considerable challenge in the pursuit of developing more comprehensive and generalizable machine learning models or databases. Specifically, to promote cross-material and general data utilization and publishing, addressing data sparsity has become an urgent priority.
For our future work, the ARIM-mdx data system plans to adopt multiple approaches. First, we aim to strengthen nationwide data collection efforts to build a more extensive materials database. Simultaneously, we will pursue integration with high-throughput automated experimental systems, enabling efficient and large-scale data generation. Furthermore, we plan to complement sparse data through the fusion of open data, academic literature, and large language models (LLMs). Through these approaches, we aim to create a more comprehensive and versatile database.
@@ VIII Related Work
Global efforts to promote open data have led to the development of several data repositories for publishing scientific data [42, 43, 44, 45]. The key difference is that our system focuses primarily on the collection and management of the raw data, rather than publishing the cleaned data. The ARIM-mdx data system provides a data platform to efficiently prepare the structured dataset for these public repositories.
The integration of large-scale materials research facilities and world-class supercomputers has been undertaken at several national research institutes such as Oak Ridge National Laboratory, Argonne National Laboratory and RIKEN [13, 14, 15, 16, 17, 18, 19, 20]. The key difference is that the ARIM-mdx data system focuses on the nationwide cross-institutional deployment.
Nationwide and cross-institutional service operations for scientific research are well-established in high-performance computing fields. The key examples are ACCESS [12] (formerly, XSEDE [10]) in the United States, EOSC [46] in the European Union, and JHPCN [11] in Japan.
@@ IX Concluding Remarks
In this paper, we introduced the ARIM-mdx data system, a nationwide data platform for materials science in Japan. The ARIM-mdx data system is based on a dedicated high-performance storage system that is tightly integrated with supercomputers, the mdx cloud system, and experimental facilities across Japan via SINET6 and the data-transfer IoT devices. The ARIM-mdx data system successfully stores, manages, and analyzes their data, facilitating more effective cross-organizational/domain research collaborations. As a result, the number of academic and industrial users approaches over 800 during the one-year trial operations. This highlights its potential as a reference design for large-scale modern materials-science data platforms.
Acknowledgment
The authors would like to thank all the contributors [^2] Contributors https://arim.mdx.jp/en/#devteam. We used mdx [47] and the HPCI/JHPCN supercomputers [^3] Supercomputers https://arim.mdx.jp/en/#supercomputer. This work was supported by MEXT ARIM Japan, “Joint Usage/Research Center for Interdisciplinary Large-scale Information Infrastructures (JHPCN)” (jh241010), MEXT “Developing a Research Data Ecosystem for the Promotion of Data-Driven Science” (2024-10) and JSPS KAKENHI (22K17899).